OpenStudy (anonymous):

What is the y-intercept of x - 2y = -6?

OpenStudy (issy14):

x-2y=-6 -2y=-x-6 y=-x/-2 -6/-2 y= 1/2x +3 m ( slope) = 1/2 y-intercept= ( 0,3)